OpenStudy (anonymous):

what is ln(ln2) - ln(ln 1)

OpenStudy (dumbcow):

ln 1 = 0 and ln 0 is undefined

OpenStudy (anonymous):

so its infinty

OpenStudy (dumbcow):

in terms of limits the answer is infinity \[\ln(\ln 2) - \lim_{x \rightarrow 1} \ln(\ln x) = \ln(\ln 2) - (-\infty) = \infty\]

OpenStudy (anonymous):

how can i find the value of p The integral from 1 to 2 of 1/(x(ln(x)^p)) dx i already Substitute u=ln(x) and you get 1/u^p

OpenStudy (anonymous):

\[\int\limits_{1}^{2}\frac{ dx }{ x (\ln x )^{p} }\]

OpenStudy (anonymous):

in to \[\int\limits_{1}^{2} \frac{ 1 }{ u ^{p} }\]

OpenStudy (anonymous):

how can i find tha value of p to make it converges

OpenStudy (dumbcow):

\[\int\limits_1^2 u^{-p} du = \frac{u^{1-p}}{1-p} = \frac{(\ln x) ^{1-p}}{1-p}\] plug in limits

OpenStudy (anonymous):

??? no idea

OpenStudy (anonymous):

why we divided by 1-p

OpenStudy (dumbcow):

power rule for integrating a polynomial \[\int\limits x^n = \frac{x^{n+1}}{n+1}\]

OpenStudy (anonymous):

how it acuatlly became u to thepower1-p

OpenStudy (anonymous):

aha

OpenStudy (anonymous):

that was stupid

OpenStudy (dumbcow):

as long as p does not equal 1, it will converge because ln(1) = 0

OpenStudy (anonymous):

so the value of p should be bigger than 1

OpenStudy (dumbcow):

p can be any number other than 1

OpenStudy (anonymous):

thank you very much you made my day <3

OpenStudy (dumbcow):

yw